At 3:00 the hour hand and the minute hand of a clock point indirections that are 90.0 degrees apart. What is the firsttime after 3:00 that the angle between the two hands has decreasedby half to 45.0 degrees?
Answer #1
For the angle between the minute hand and the hour hand todecrease by 45o, the minute hand must rotate through45o relative to the hour hand. Both are moving in theclockwise direction - the min hand faster.
The angular speeds are, ω1 = 2π rad/ 60min, for the min hand and ω2 = 2π rad/12 hour,for the hour hand.
(each takes 60 min and 12 hour respectively to make onecomplete rotation or 2π radians.).
Rel speed of the min hand with respect to the hourhand,
            ω = ω1 - ω2 =(2π/60)(1 - 1/12) = 11π/360 rad/min.
At this speed it must cover, θ=  45o = π/4 rad.
Using     θ = ωt
Time taken,
            t = θ/ω = (π/4)/11π/360 = 90/11
               = 8.1818 min = 8.18min.
This is at time3:08:11 (8 min and 11 s past 3:00).
Answer #2
For theangle between the minute hand and the hour hand to decrease by45o, the minute hand must rotate through 45orelative to the hour hand. Both are moving in the clockwisedirection - the min hand faster.
Theangular speeds are, ω1 = 2π rad/ 60 min, forthe min hand and ω2 = 2π rad/12 hour, for thehour hand.
(eachtakes 60 min and 12 hour respectively to make one complete rotationor 2π radians.).
Relspeed of the min hand with respect to the hourhand,
            ω = ω1 - ω2 =(2π/60)(1 - 1/12) = 11π/360 rad/min.
At thisspeed it must cover, θ =  45o = π/4rad.
Using    θ = ωt
Timetaken,
            t = θ/ω = (π/4)/11π/360 =90/11
               = 8.1818 min = 8.18min.
This isat time 3:08:11 (8min and 11 s past 3:00).

I do not understand how you got the relative speed of the minutehand with respect to the hour hand. How do you know to useω= ω1-ω2?